# Preserving Membership Privacy in Low-Rank Adaptation for Latent Diffusion Models

### v2

@@ Abstract
Low-rank adaptation (LoRA) is an efficient strategy for adapting latent diffusion models (LDMs) on a private dataset to generate specific images by minimizing the adaptation loss. However, the LoRA-adapted LDMs are vulnerable to membership inference (MI) attacks that can judge whether a particular data point belongs to the private dataset, thus leading to the privacy leakage. To defend against MI attacks, we first propose a straightforward solution: Membership-Privacy-preserving LoRA (MP-LoRA). MP-LoRA is formulated as a min-max optimization problem where a proxy attack model is trained by maximizing its MI gain while the LDM is adapted by minimizing the sum of the adaptation loss and the MI gain of the proxy attack model. However, we empirically find that MP-LoRA has the issue of unstable optimization, and theoretically analyze that the potential reason is the unconstrained local smoothness, which impedes the privacy-preserving adaptation. To mitigate this issue, we further propose a Stable Membership-Privacy-preserving LoRA (SMP-LoRA) that adapts the LDM by minimizing the ratio of the adaptation loss to the MI gain. Besides, we theoretically prove that the local smoothness of SMP-LoRA can be constrained by the gradient norm, leading to improved convergence. Our experimental results corroborate that SMP-LoRA can indeed defend against MI attacks and generate high-quality images. Our code is available at GitHub.
$\ast$\$$\ast$\$footnotetext: Equal contribution.$\dagger$\$$\dagger$\$footnotetext: Correspondence to: Jingfeng Zhang <jingfeng.zhang@auckland.ac.nz>.
@@ 1 Introduction
Generative diffusion models (Ho et al., 2020; Song et al., 2020) are leading a revolution in AI-generated content, renowned for their unique generation process and fine-grained image synthesis capabilities. Notably, the Latent Diffusion Model (LDM) (Rombach et al., 2022; Podell et al., 2023) stands out by executing the diffusion process in latent space, enhancing computational efficiency without compromising image quality. Thus, LDMs can be efficiently adapted to generate previously unseen contents or styles (Meng et al., 2021; Gal et al., 2022; Ruiz et al., 2023; Zhang et al., 2023), thereby catalyzing a surge across multiple fields, such as facial generation (Huang et al., 2023; Xu et al., 2024) and medicine (Kazerouni et al., 2022; Shavlokhova et al., 2023).
@@
Among various adaptation methods, Low-Rank Adaptation (LoRA) (Hu et al., 2021) is the superior strategy for adapting LDMs by significantly reducing computational resources while ensuring commendable performance with great flexibility. Compared to the full fine-tuning method which fine-tunes all parameters, LoRA optimizes the much smaller low-rank matrices, making the training more efficient and lowering the hardware requirements for adapting LDMs (Hu et al., 2021). By performing the low-rank decomposition of the transformer structure within the LDM, LoRA offers performance comparable to fine-tuning all LDM parameters (Cuenca and Paul, 2023). Moreover, LoRA allows flexible sharing of a pre-trained LDM to build numerous small LoRA modules for various tasks.
@@
However, recent studies (Pang and Wang, 2023; Pang et al., 2023; Dubiński et al., 2024) have pointed out that adapted LDMs are facing the severe risk of privacy leakage. The leakage primarily manifests in the vulnerability to Membership Inference (MI) attacks (Shokri et al., 2017), which utilize the model’s loss of a data point to differentiate whether it is a member of the training dataset or not. As shown in Figure 1(d), the LoRA-adapted LDM (red circle marker) exhibits an incredibly high Attack Success Rate (ASR) of 74.1%.
(a) Training loss.
(b) Gradient scale.
(c) Generated results.
(d) Performance.
@@
Figure 1: Figure 1(a) shows the trajectory of the training loss during the adaptation process via LoRA, MP-LoRA, and SMP-LoRA on the Pokemon dataset. Figure 1(b) displays the mean and standard deviation of the gradient norms and Hessian norms for MP-LoRA and SMP-LoRA throughout the training iterations. It also presents the Pearson correlation coefficients (PCC) and p-values assessing their correlation. Note that each epoch contains 433 training iterations. Figures 1(c) and 1(d) demonstrate the generated images and a comparison of evaluation metrics including FID Score and MI attack success rate (ASR). MP-LoRA preserves membership privacy at the cost of diminished image generation capability. In contrast, SMP-LoRA effectively preserves membership privacy while maintaining the quality of the generated image, demonstrating its effectiveness in defending against MI attacks without significant loss of functionality. Extensive generated images are visualized in Appendix 6.9.
@@
However, the vanilla MP-LoRA encounters an issue of effective optimization of the training loss, as evidenced in the orange dashed line of Figure 1(a). We theoretically find that during MP-LoRA, the local smoothness, quantified by the Hessian norm (the norm of the Hessian matrix) (Bubeck et al., 2015), is independent of and not bounded by the gradient norm (see Proposition 1 for detials). This independence hinders the privacy-preserving adaptation of the MP-LoRA (Zhang et al., 2019), thus impeding optimizing the training loss. Besides, we empirically show that the correlation between the Hessian norm and the gradient norm during MP-LoRA is insignificant. This is manifested by the Pearson Correlation Coefficient (PCC) of 0.043 and the p-value above 0.05, as shown in the upper panel of Figure 1(b), which corroborates our theoretical analyses.
@@
To evaluate the performance of the SMP-LoRA, we conducted adapting experiments using the Stable Diffusion v1.5 (CompVis, 2022) on the Pokemon (Pinkney, 2022) and CelebA (Liu et al., 2015) datasets, respectively. Figure 1(d) shows that, although MP-LoRA (orange square marker) lowers the ASR to near-random levels, it significantly degrades the image generation capability of LoRA, as evidenced by a high FID score of 3.513 and the poor visual quality in Figure 1(c). In contrast, the SMP-LoRA (blue pentagon marker) effectively preserves membership privacy without sacrificing generated image quality significantly, as evidenced by its FID score of 0.274 and ASR of 50.3%.
@@ 2 Background and Preliminary
2.1 Diffusion Model
Diffusion models (DMs) (Ho et al., 2020; Song et al., 2020) have shown remarkable performance in image synthesis. Compared with other generative models such as GAN (Goodfellow et al., 2014), DMs can mitigate the problems of the training instability and the model collapse (Rombach et al., 2022), while achieving state-of-the-art results in numerous benchmarks (Dhariwal and Nichol, 2021). Among various implementations of DMs, the Latent Diffusion Model (LDM) is renowned for generating high-quality images with limited computational resources, thereby widely utilized across many applications. Fundamentally, LDM is a Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) built in the latent space, effectively reducing computational demand while enabling high-quality and flexible image generation (Rombach et al., 2022). Therefore, LDMs have been widely utilized for adaptation (Gal et al., 2022; Hu et al., 2021; Ruiz et al., 2023), capable of delivering high-performance models even when adapted on small datasets.
@@ 2.1 Diffusion Model
where $\tau_{\phi}$ refers to the pre-trained text encoder from CLIP (Radford et al., 2021).
2.2 Low-Rank Adaptation (LoRA)
To unleash the power of large pretrained models, many adaptation methods have emerged. In particular, Low-Rank Adaptation (LoRA) (Hu et al., 2021) provides an efficient and effective solution by freezing the pre-trained model weights and introducing the trainable low-rank counterparts, significantly reducing the number of trainable parameters and memory usage during the LoRA adaptation process. Therefore, LoRA not only lessens the demand for computational resources but also allows for the construction of multiple lightweight portable low-rank matrices on the same pre-trained LDM, addressing various downstream tasks (Cuenca and Paul, 2023; Huggingface, 2023).
@@ 2.2 Low-Rank Adaptation (LoRA)
2.3 Membership Inference Attack
Membership Inference (MI) attack (Shokri et al., 2017) aims to determine whether a particular data point is part of a model’s training set. Recent studies (Wu et al., 2022; Hu and Pang, 2023; Matsumoto et al., 2023; Carlini et al., 2023; Pang et al., 2023; Duan et al., 2023; Kong et al., 2023) have shown that DMs are particularly vulnerable to MI attacks, thus undergoing high risks of privacy leakage.
As a pioneering work to explore MI attacks on DMs, Wu et al. (2022) leveraged the intuition that when replicating training images, DMs tend to produce images with a higher degree of fidelity and greater consistency with the corresponding textual captions. They conducted black-box MI attacks on text-to-image DMs at both pixel and semantic levels. Compared with the pixel level, the semantic-level attack achieved better attack performance. Specifically, the semantic-level attack utilized the L2 distance between the embeddings of a given image and its corresponding caption-generated image to make inferences.
For white-box attacks, Hu and Pang (2023) and Matsumoto et al. (2023) employed a threshold-based MI attack using the model loss at specific diffusion steps, as defined in Equation (1), and yielded significant attack performance. The threshold-based MI attack identifies training data by assessing whether the loss value for a given data point falls below a certain threshold. Meanwhile, Pang et al. (2023) utilized the model’s gradient to train an attack model for inference, also achieving remarkable attack performance.
Despite adapted LDMs facing a severe risk of privacy leakage (Pang and Wang, 2023; Pang et al., 2023; Dubiński et al., 2024), no effective strategy has been proposed to protect membership privacy so far. To mitigate the privacy risk, our paper takes the first step by developing a defensive adaptation against MI attacks for DMs.
@@ 3 Membership-Privacy-Preserving LoRA
3.1 A Vanilla Solution: MP-LoRA
@@
In MI attack, the conflicting objectives of defenders and adversaries can be modelled as a privacy game (Shokri et al., 2012; Manshaei et al., 2013; Alvim et al., 2017). Adversaries can adjust their attack models to maximize MI gain against the target model, which requires that the defense can anticipate and withstand the strongest inference attacks. Consequently, the defender’s goal is to enhance the preservation of membership privacy in worst-case scenarios where the adversary achieves the maximum MI gain while maintaining the model performance. Inspired by Nasr et al. (2018), we propose MP-LoRA to defend against MI attacks which is formulated as a min-max optimization problem as follows:
@@
$G\left(h_{\omega},\mathcal{D}_{\mathrm{aux}},f_{\bar{\theta}+\mathbf{B}\mathbf{A}}\right)=$ $\frac{1}{2\left|\mathcal{D}^{\mathrm{m}}_{\mathrm{aux}}\right|}\sum_{(x,y)\in\mathcal{D}^{\mathrm{m}}_{\mathrm{aux}}}\log{\left(h_{\omega}\left(\ell_{\mathrm{ada}}\left(x,y;f_{\bar{\theta}+\mathbf{B}\mathbf{A}}\right)\right)\right)}$
@@
MP-LoRA is realized by one step of inner maximization to obtain a power proxy attack model by maximizing the MI gain in Equation (3.1) and one step of outer minimization to update $\mathbf{A}$ and $\mathbf{B}$ by minimizing the training loss in Equation (6). The algorithm of MP-LoRA is shown in Algorithm 2 (Appendix 6.1).
3.2 Unstable Issue of MP-LoRA
@@
$\|\frac{\partial^{2}\mathcal{L}_{\mathrm{PL}}}{\partial\mathbf{B}\mathbf{A}^{2}}\|\leq L_{0}+L_{1}\|\frac{\partial\mathcal{L}_{\mathrm{PL}}}{\partial\mathbf{B}\mathbf{A}}\|,~{}\mathrm{where}~{}~{}L_{0}=\frac{\partial^{2}\mathcal{L}_{\mathrm{ada}}}{\partial\mathbf{B}\mathbf{A}^{2}}+\lambda\frac{\partial^{2}G}{\partial\mathbf{B}\mathbf{A}^{2}},~{}L_{1}=0,$ (8)
in which $\mathcal{L}_{\mathrm{ada}}$ represents the adaptation loss and $G$ represents the MI gain. The detailed derivation is presented in Appendix 6.2. The value of $L_{1}$ being zero indicates that the Hessian norm is independent of and not bounded by the gradient norm, suggesting that the local smoothness is unconstrained. ∎
Next, we provide empirical evidence to support our theoretical analyses. We tracked the gradient norm and the Hessian norm of the training loss at each training iteration, and calculated their Pearson Correlation coefficient (PCC) and p-value as shown in Figure 1(b). The details for calculating the gradient norm and the Hessian norm can be found in Appendix 6.3. In Figure 1(b), the low PPC of 0.043 for MP-LoRA suggests a very weak correlation between the Hessian norm and the gradient norm. Additionally, with the p-value of 0.052, there is insufficient evidence to reject the hypothesis of no correlation. This indicates that the Hessian norm is unbounded, implying that the local smoothness, quantified by the Hessian norm (Bubeck et al., 2015), is unconstrained. Such unconstrained local smoothness leads to the unstable optimization issue in MP-LoRA, and even to the failure of adaptation, as evidenced in the orange dashed line of Figure 1(a) and the poor visual quality of the generated images in Figure 1(c).
3.3 Stabilizing MP-LoRA
@@
$\|\frac{\partial^{2}\mathcal{L}_{\mathrm{SPL}}}{\partial\mathbf{B}\mathbf{A}^{2}}\|\leq L^{\prime}_{0}+L^{\prime}_{1}\|\frac{\partial\mathcal{L}_{\mathrm{SPL}}}{\partial\mathbf{B}\mathbf{A}}\|,~{}\mathrm{where}~{}~{}\mu=\frac{\partial\mathcal{L}_{\mathrm{ada}}}{\partial\mathbf{B}\mathbf{A}},~{}\nu=\lambda\frac{\partial G}{\partial\mathbf{B}\mathbf{A}},$ (11)
@@
$L^{\prime}_{0}=\frac{1}{1-\lambda G+\delta}\cdot\frac{\partial^{2}\mathcal{L}_{\mathrm{ada}}}{\partial\mathbf{B}\mathbf{A}^{2}}+\frac{\lambda\mathcal{L}_{\mathrm{ada}}}{(1-\lambda G+\delta)^{2}}\cdot\frac{\partial G^{2}}{\partial\mathbf{B}\mathbf{A}^{2}},~{}L^{\prime}_{1}=\frac{2\nu}{1-\lambda G+\delta}.$
@@
Please refer to Appendix 6.2 for detailed derivation. The value of $L^{\prime}_{1}$ being greater than zero indicates that the Hessian norm is positively correlated with and upper bounded by the gradient norm, suggesting that the gradient norm constrains the local smoothness during adaptation. ∎
Notably, SMP-LoRA also exhibits lower mean and standard deviation of the gradient norm compared to MP-LoRA. We provide further empirical analysis in Appendix 6.4, showing that SMP-LoRA can implicitly rescale the gradient during adaptation by introducing the factors $\frac{1}{1-\lambda G+\delta}$ and $\frac{\mathcal{L}_{\mathrm{ada}}}{\left(1-\lambda G+\delta\right)^{2}}$, thereby leading to more stable gradient and controlled gradient scale compared to MP-LoRA.
Algorithm 1 SMP-LoRA
Output: a membership-privacy-preserving LoRA for LDMs.
Perform low-rank decomposition on $f_{\theta}$ to obtain $f_{\bar{\theta}+\mathbf{B}\mathbf{A}}$ ($\mathbf{B}$ and $\mathbf{A}$ are trainable LoRA modules.)
for each epoch do
     for each training iteration do
         Sample batches $S^{\mathrm{m}}$ and $S^{\mathrm{nm}}$ from $\mathcal{D}^{\mathrm{m}}_{\mathrm{aux}}$ and $\mathcal{D}^{\mathrm{nm}}_{\mathrm{aux}}$, respectively
         Calculate the MI gain $G^{*}=G(h_{\omega},S^{\mathrm{m}}\cup S^{\mathrm{nm}},f_{\bar{\theta}+\mathbf{B}\mathbf{A}})$
         Update the parameters $\omega\leftarrow\omega+\eta_{1}\cdot\nabla_{\omega}G^{*}$
         Sample a fresh batch from $\mathcal{D}_{\mathrm{tr}}$
         Calculate the training loss $\mathcal{L}^{*}=\mathcal{L}_{\mathrm{SPL}}$
         Update parameters $\mathbf{A}\leftarrow\mathbf{A}-\eta_{2}\cdot\nabla_{\mathbf{A}}\mathcal{L}^{*}$ and $\mathbf{B}\leftarrow\mathbf{B}-\eta_{2}\cdot\nabla_{\mathbf{B}}\mathcal{L}^{*}$, respectively
     end for
end for
4 Experiments
In this section, we first evaluate the performance of LoRA, MP-LoRA, and SMP-LoRA in terms of image generation capability and effectiveness in defending against MI attacks. Then, we conduct ablation studies on the important hyperparameters and further extend our membership-privacy-preserving method to full fine-tuning and DreamBooth (Ruiz et al., 2023) methods. Subsequently, we compare SMP-LoRA with traditional techniques for stabilizing the gradient and evaluate the effectiveness of SMP-LoRA in defending against MI attacks in different settings. For details on the experimental setup, please refer to Appendix 6.5.
Dataset.
In our experiment, we utilized two datasets: Pokemon (Pinkney, 2022) and CelebA (Liu et al., 2015). From CelebA, we created four subsets: CelebA__\__Small and CelebA__\__Large, both balanced with equal image contributions per individual, and CelebA__\__Gender and CelebA__\__Varying, imbalanced with a $7:3$ gender ratio and varied image contributions per individual, respectively. For more details, please refer to Appendix 6.5.
Model hyperparameters.
We utilized the official pre-trained Stable Diffusion v1.5 (CompVis, 2022) for LDMs to build the LoRA module, with specific model hyperparameters detailed in Table 5 in Appendix 6.5. We employed a 3-layer MLP as the proxy attack model $h_{\omega}$ and a structurally similar new attack model $h^{\prime}$ to evaluate the effectiveness of adapted LDMs in defending against MI attacks.
Evaluation metrics.
We employed Attack Success Rate (ASR) (Choquette-Choo et al., 2021), Area Under the ROC Curve (AUC), and True Positive Rate (TPR) to evaluate the effectiveness of MP-LoRA and SMP-LoRA in defending against MI attacks. Lower values for ASR and TPR indicate a more effective defense against MI attacks. Consistent with many prior studies (Chen et al., 2022; Ye et al., 2022; Duan et al., 2023; Dubiński et al., 2024), we default that the TPR measures the capability of the attack model to identify samples as members correctly. Consequently, an AUC value closer to 0.5, i.e., a smaller value of $\frac{\left|\mathrm{AUC}-0.5\right|}{0.5}$, represents a stronger defense capability, as MI attacks involve determining both membership and non-membership. For assessing the image generation capability of the adapted LDMs, we utilized the Fréchet Inception Distance (FID) (Heusel et al., 2017) and the Kernel Inception Distance (KID) (Bińkowski et al., 2018), with lower values denoting better image quality. Please refer to Appendix 6.5 for detailed explanations.

### v3

# Privacy-Preserving Low-Rank Adaptation against Membership Inference Attacks for Latent Diffusion Models

@@ Abstract
Low-rank adaptation (LoRA) is an efficient strategy for adapting latent diffusion models (LDMs) on a private dataset to generate specific images by minimizing the adaptation loss. However, the LoRA-adapted LDMs are vulnerable to membership inference (MI) attacks that can judge whether a particular data point belongs to the private dataset, thus leading to the privacy leakage. To defend against MI attacks, we first propose a straightforward solution: Membership-Privacy-preserving LoRA (MP-LoRA). MP-LoRA is formulated as a min-max optimization problem where a proxy attack model is trained by maximizing its MI gain while the LDM is adapted by minimizing the sum of the adaptation loss and the MI gain of the proxy attack model. However, we empirically find that MP-LoRA has the issue of unstable optimization, and theoretically analyze that the potential reason is the unconstrained local smoothness, which impedes the privacy-preserving adaptation. To mitigate this issue, we further propose a Stable Membership-Privacy-preserving LoRA (SMP-LoRA) that adapts the LDM by minimizing the ratio of the adaptation loss to the MI gain. Besides, we theoretically prove that the local smoothness of SMP-LoRA can be constrained by the gradient norm, leading to improved convergence. Our experimental results corroborate that SMP-LoRA can indeed defend against MI attacks and generate high-quality images.
Codehttps://github.com/WilliamLUO0/StablePrivateLoRA
@@ 1 Introduction
Generative diffusion models (Ho, Jain, and Abbeel 2020; Song et al. 2021) are leading a revolution in AI-generated content, renowned for their unique generation process and fine-grained image synthesis capabilities. Notably, the Latent Diffusion Model (LDM) (Rombach et al. 2022; Podell et al. 2024) stands out by executing the diffusion process in latent space, enhancing computational efficiency without compromising image quality. Thus, LDMs can be efficiently adapted to generate previously unseen contents or styles (Meng et al. 2022; Gal et al. 2023; Ruiz et al. 2023; Zhang, Rao, and Agrawala 2023), thereby catalyzing a surge across multiple fields, such as facial generation (Huang et al. 2023; Xu et al. 2024) and medicine (Kazerouni et al. 2022; Shavlokhova et al. 2023).
@@
Among various adaptation methods, Low-Rank Adaptation (LoRA) (Hu et al. 2022) is the superior strategy for adapting LDMs by significantly reducing computational resources while ensuring commendable performance with great flexibility. Compared to the full fine-tuning method which fine-tunes all parameters, LoRA optimizes the much smaller low-rank matrices, making the training more efficient and lowering the hardware requirements for adapting LDMs (Hu et al. 2022). By performing the low-rank decomposition of the transformer structure within the LDM, LoRA offers performance comparable to fine-tuning all LDM parameters (Cuenca and Paul 2023). Moreover, LoRA allows flexible sharing of a pre-trained LDM to build numerous small LoRA modules for various tasks.
@@
However, recent studies (Pang and Wang 2023; Pang et al. 2023; Dubiński et al. 2024) have pointed out that adapted LDMs are facing the severe risk of privacy leakage. The leakage primarily manifests in the vulnerability to Membership Inference (MI) attacks (Shokri et al. 2017), which utilize the model’s loss of a data point to differentiate whether it is a member of the training dataset or not. As shown in Figure 1(d), the LoRA-adapted LDM (red circle marker) exhibits an incredibly high Attack Success Rate (ASR) of 82.27%.
(a) Training loss
(b) Gradient scale
(c) Generated results
(d) Performance
@@
Figure 1: Figure 1(a) shows the trajectory of the training loss during the adaptation process via LoRA, MP-LoRA, and SMP-LoRA on the Pokemon dataset. Figure 1(b) displays the mean and standard deviation of the gradient norms and Hessian norms for MP-LoRA and SMP-LoRA throughout the training iterations. It also presents the Pearson correlation coefficients (PCC) and p-values assessing their correlation. Note that each epoch contains 433 training iterations. Figures 1(c) and 1(d) demonstrate the generated images and a comparison of evaluation metrics including FID Score and MI attack success rate (ASR). MP-LoRA preserves membership privacy but compromises image generation capability. In contrast, SMP-LoRA effectively preserves membership privacy while maintaining the quality of the generated image, demonstrating its effectiveness in defending against MI attacks without significant loss of functionality. Extensive generated images are visualized in Appendix H.
@@
However, the vanilla MP-LoRA encounters an issue of effective optimization of the training loss, as evidenced in the orange dashed line of Figure 1(a). We theoretically find that during MP-LoRA, the local smoothness, quantified by the Hessian norm (the norm of the Hessian matrix) (Bubeck et al. 2015), is independent of and not bounded by the gradient norm (see Proposition 1 for details). This independence hinders the privacy-preserving adaptation of the MP-LoRA (Zhang et al. 2019), thus impeding optimizing the training loss. Besides, we empirically show that the correlation between the Hessian norm and the gradient norm during MP-LoRA is insignificant. This is manifested by the Pearson Correlation Coefficient (PCC) of 0.043 and the p-value above 0.05, as shown in the upper panel of Figure 1(b), which corroborates our theoretical analyses.
@@
To evaluate the performance of the SMP-LoRA, we conducted adapting experiments using the Stable Diffusion v1.5 (CompVis 2022) on the Pokemon (Pinkney 2022) and CelebA (Liu et al. 2015) datasets, respectively. Figure 1(d) shows that, although MP-LoRA (orange square marker) lowers the ASR to near-random levels, it significantly degrades the image generation capability of LoRA, as evidenced by a high FID score of 2.10 and the poor visual quality in Figure 1(c). In contrast, the SMP-LoRA (blue pentagon marker) effectively preserves membership privacy without sacrificing generated image quality significantly, as evidenced by its FID score of 0.32 and ASR of 51.97%.
@@ 2 Background and Preliminary
Diffusion Model
Diffusion models (DMs) (Ho, Jain, and Abbeel 2020; Song et al. 2021) have shown remarkable performance in image synthesis. Compared with other generative models such as GAN (Goodfellow et al. 2014), DMs can mitigate the problems of the training instability and the model collapse (Rombach et al. 2022), while achieving state-of-the-art results in numerous benchmarks (Dhariwal and Nichol 2021). Among various implementations of DMs, the Latent Diffusion Model (LDM) is renowned for generating high-quality images with limited computational resources, thereby widely utilized across many applications. Fundamentally, LDM is a Denoising Diffusion Probabilistic Model (DDPM) (Ho, Jain, and Abbeel 2020) built in the latent space, effectively reducing computational demand while enabling high-quality and flexible image generation (Rombach et al. 2022). Therefore, LDMs have been widely utilized for adaptation (Gal et al. 2023; Hu et al. 2022; Ruiz et al. 2023), capable of delivering high-performance models even when adapted on small-scale datasets.
@@
where $\tau_{\phi}$ refers to the pre-trained text encoder from CLIP (Radford et al. 2021).
Low-Rank Adaptation (LoRA)
To unleash the power of large pretrained models, many adaptation methods have emerged. In particular, Low-Rank Adaptation (LoRA) (Hu et al. 2022) provides an efficient and effective solution by freezing the pre-trained model weights and introducing the trainable low-rank counterparts, significantly reducing the number of trainable parameters and memory usage during the LoRA adaptation process. Therefore, LoRA not only lessens the demand for computational resources but also allows for the construction of multiple lightweight portable low-rank matrices on the same pre-trained LDM, addressing various downstream tasks (Cuenca and Paul 2023; Huggingface 2023).
@@
Membership Inference Attack
Membership Inference (MI) attack (Shokri et al. 2017) aims to determine whether a particular data point is part of a model’s training set. Recent studies (Carlini et al. 2023; Duan et al. 2023) have shown that DMs are particularly vulnerable to MI attacks, thus undergoing high risks of privacy leakage. MI attacks on diffusion models can be categorized based on the adversary’s capabilities into white-box (Hu and Pang 2023; Matsumoto, Miura, and Yanai 2023; Pang et al. 2023), gray-box (Duan et al. 2023; Kong et al. 2024; Fu et al. 2023), and black-box (Wu et al. 2022; Matsumoto, Miura, and Yanai 2023; Pang and Wang 2023; Zhang et al. 2024) attacks.
In the black-box setting, Wu et al. (2022) were the first to explore MI attacks on DMs and achieved the highest success rate among the black-box attacks mentioned above. They noted that DMs, when replicating training images, consistently produce outputs with higher fidelity and greater alignment with textual captions, indicating significant behavioural differences. Therefore, they utilized the L2 distance between the embeddings of a given image and its corresponding caption-generated image to infer membership.
In the white-box setting, the gradient-based MI attack developed by Pang et al. (2023) stands as the most effective method for DMs in terms of attack performance. They leveraged the model’s gradient to train an attack model for inference. Additionally, Hu and Pang (2023) and Matsumoto, Miura, and Yanai (2023) employed a threshold-based MI attack by analyzing model loss at specific diffusion steps, which we refer to as the loss-based MI attack, also yielding significant attack performance. In contrast, the black-box and gray-box MI attacks previously mentioned, which lack access to internal model information, achieve lower success rates compared to the white-box loss-based and gradient-based MI attacks (Pang et al. 2023).
Conventional techniques to defend against MI attacks include differential privacy (Dwork 2008; Abadi et al. 2016), min-max membership privacy game (Nasr, Shokri, and Houmansadr 2018), data augmentation (DeVries and Taylor 2017; Cubuk et al. 2020), early stopping, etc. To the best of our knowledge, these techniques have not been systematically applied or comprehensively evaluated in DMs. Notably, some studies (Dockhorn et al. 2023; Ghalebikesabi et al. 2023; Lyu et al. 2023) have applied differential privacy to DMs to achieve a better privacy-accuracy trade-off, yet they have not evaluated its effectiveness against MI attacks in DMs. Therefore, our paper takes the first step by developing a defensive adaptation based on the min-max membership privacy game to defend against MI attacks in DMs.
@@ 3 Membership-Privacy-Preserving LoRA
A Vanilla Solution: MP-LoRA
@@
In MI attack, the conflicting objectives of defenders and adversaries can be modelled as a privacy game (Shokri et al. 2012; Manshaei et al. 2013; Alvim et al. 2017). Adversaries can adjust their attack models to maximize MI gain against the target model, which requires that the defense can anticipate and withstand the strongest inference attacks. Consequently, the defender’s goal is to enhance the preservation of membership privacy in worst-case scenarios where the adversary achieves the maximum MI gain while maintaining the model performance. Inspired by Nasr, Shokri, and Houmansadr (2018), we propose MP-LoRA to defend against MI attacks which is formulated as a min-max optimization problem as follows:
@@
$G\left(h_{\omega},\mathcal{D}_{\mathrm{aux}},f_{\bar{\theta}+\mathbf{B}\mathbf{A}}\right)=$
$\frac{1}{2\left|\mathcal{D}^{\mathrm{m}}_{\mathrm{aux}}\right|}\sum_{(x,y)\in\mathcal{D}^{\mathrm{m}}_{\mathrm{aux}}}\log{\left(h_{\omega}\left(\ell_{\mathrm{ada}}\left(x,y;f_{\bar{\theta}+\mathbf{B}\mathbf{A}}\right)\right)\right)}$
@@
MP-LoRA is realized by one step of inner maximization to obtain a power proxy attack model by maximizing the MI gain in Equation (3) and one step of outer minimization to update $\mathbf{A}$ and $\mathbf{B}$ by minimizing the training loss in Equation (6). The algorithm of MP-LoRA is shown in Algorithm 2 (Appendix A).
Unstable Issue of MP-LoRA
@@
$\|\frac{\partial^{2}\mathcal{L}_{\mathrm{PL}}}{\partial\mathbf{B}\mathbf{A}^{2}}\|\leq L_{0}+L_{1}\|\frac{\partial\mathcal{L}_{\mathrm{PL}}}{\partial\mathbf{B}\mathbf{A}}\|,$
$\mathrm{where}~{}~{}L_{0}=\|\frac{\partial^{2}\mathcal{L}_{\mathrm{ada}}}{\partial\mathbf{B}\mathbf{A}^{2}}\|+\lambda\|\frac{\partial^{2}G}{\partial\mathbf{B}\mathbf{A}^{2}}\|,~{}L_{1}=0,$ (8)
in which $\mathcal{L}_{\mathrm{ada}}$ represents the adaptation loss and $G$ represents the MI gain. The detailed derivation is presented in Appendix B. The value of $L_{1}$ being zero indicates that the Hessian norm is independent of and not bounded by the gradient norm, suggesting that the local smoothness is unconstrained. ∎
Next, we provide empirical evidence to support our theoretical analyses. We tracked the gradient norm and the Hessian norm of the training loss at each training iteration, and calculated their Pearson Correlation coefficient (PCC) and p-value as shown in Figure 1(b). The details for calculating the gradient norm and the Hessian norm can be found in Appendix C. In Figure 1(b), the low PPC of 0.043 for MP-LoRA suggests a very weak correlation between the Hessian norm and the gradient norm. Additionally, with the p-value of 0.052, there is insufficient evidence to reject the hypothesis of no correlation. This indicates that the Hessian norm is unbounded, implying that the local smoothness, quantified by the Hessian norm (Bubeck et al. 2015), is unconstrained. Such unconstrained local smoothness leads to the unstable optimization issue in MP-LoRA, and even to the failure of adaptation, as evidenced in the orange dashed line of Figure 1(a) and the poor visual quality of the generated images in Figure 1(c).
Stabilizing MP-LoRA
@@
Algorithm 1 Stable Membership-Privacy-preserving LoRA
Output: a SMP-LoRA for LDMs
1:  Perform low-rank decomposition on $f_{\theta}$ to obtain $f_{\bar{\theta}+\mathbf{B}\mathbf{A}}$ ($\mathbf{B}$ and $\mathbf{A}$ are trainable LoRA modules)
2:  for each epoch do
3:     for each training iteration do
4:        Sample batches $S^{\mathrm{m}}$ and $S^{\mathrm{nm}}$ from $\mathcal{D}^{\mathrm{m}}_{\mathrm{aux}}$ and $\mathcal{D}^{\mathrm{nm}}_{\mathrm{aux}}$
5:        Calculate the MI gain $G^{*}$ on $S^{\mathrm{m}}\cup S^{\mathrm{nm}}$
6:        Update the parameters $\omega\leftarrow\omega+\eta_{1}\cdot\nabla_{\omega}G^{*}$.
7:        Sample a fresh batch from $\mathcal{D}_{\mathrm{tr}}$
8:        Calculate the training loss $\mathcal{L}^{*}=\mathcal{L}_{\mathrm{SPL}}$
9:        Update parameters $\mathbf{A}\leftarrow\mathbf{A}-\eta_{2}\cdot\nabla_{\mathbf{A}}\mathcal{L}^{*}$ and $\mathbf{B}\leftarrow\mathbf{B}-\eta_{2}\cdot\nabla_{\mathbf{B}}\mathcal{L}^{*}$, respectively
10:     end for
11:  end for
@@
$\|\frac{\partial^{2}\mathcal{L}_{\mathrm{SPL}}}{\partial\mathbf{B}\mathbf{A}^{2}}\|\leq L^{\prime}_{0}+L^{\prime}_{1}\|\frac{\partial\mathcal{L}_{\mathrm{SPL}}}{\partial\mathbf{B}\mathbf{A}}\|,$
$\mathrm{where}~{}~{}\mu=\frac{\partial\mathcal{L}_{\mathrm{ada}}}{\partial\mathbf{B}\mathbf{A}},~{}\nu=\lambda\frac{\partial G}{\partial\mathbf{B}\mathbf{A}},$
@@
$L^{\prime}_{0}=\frac{1}{1-\lambda G+\delta}\cdot\|\frac{\partial^{2}\mathcal{L}_{\mathrm{ada}}}{\partial\mathbf{B}\mathbf{A}^{2}}\|+\frac{\lambda\mathcal{L}_{\mathrm{ada}}}{(1-\lambda G+\delta)^{2}}\cdot\|\frac{\partial G^{2}}{\partial\mathbf{B}\mathbf{A}^{2}}\|,$
$L^{\prime}_{1}=\frac{2\|\nu\|}{1-\lambda G+\delta}.$ (11)
@@
Please refer to Appendix B for detailed derivation. The value of $L^{\prime}_{1}$ being greater than zero indicates that the Hessian norm is positively correlated with and upper bounded by the gradient norm, suggesting that the gradient norm constrains the local smoothness during adaptation. ∎